Solve the radical equation.
6 SQRT of x = x+5

Answer by fractalier(2101)

Okay from 6sqrt(x) = x + 5, let us square both sides...
36x = x^2 + 10x + 25
Subtract the 36x, factor, solve and check...
x^2 - 26x + 25 = 0
(x - 25)(x - 1) = 0
x = 25 or x = 1
They both check out okay...two solutions here...